### v3

# On a sequence of Kimberling and its relationship to the Tribonacci word

Abstract
In 2017, Clark Kimberling defined an interesting sequence ${\bf B}=0100101100\cdots$ of 0’s and $1$’s by certain inflation rules, and he made a number of conjectures about this sequence and some related ones. In this note we prove his conjectures using, in part, the Walnut theorem-prover. We show how his word is related to the infinite Tribonacci word, and we determine both the subword complexity and critical exponent of $\bf B$.
1 Introduction
In June 2017, Clark Kimberling defined sequence A288462 in the OEIS [17] as follows: it is the infinite fixed point of the inflation rules $00\rightarrow 0101$, $1\rightarrow 10$, starting with $00$. In this note we prove his conjectures using, in part, the Walnut theorem-prover. We also show how his word is related to the infinite Tribonacci word, and we determine both the subword complexity and critical exponent of $\bf B$.
It may seem odd to devote an entire paper to a particular sequence, but the methods we use are widely applicable, and so the paper may serve as a primer on how to attack such a sequence using a combination of combinatorial and automata-theoretic techniques.
As stated this description is perhaps slightly vague, so here is some elaboration. We start with $B_{0}=00$. To find $B_{i+1}$ from $B_{i}$, we do the following: we factor $B_{i}$ into maximal blocks of the form $00$, 0, and $1$; here maximal means we cannot extend a block further to the right or left. Then $B_{i+1}$ is the result of applying the inflation rules $0\rightarrow 0$, $1\rightarrow 10$, and $00\rightarrow 0101$ to $B_{i}$.
For example, here are the first few iterates:
$B_{0}$ $=00$
$B_{1}$ $=0101$
$B_{2}$ $=010010$
$B_{3}$ $=0100101100$
$B_{4}$ $=010010110010100101$
$B_{5}$ $=01001011001010010110010010110010.$
It is not hard to see that $B_{i+1}$ is a prefix of $B_{i}$ for $i\geq 2$, and so there is a unique infinite word ${\bf B}=a_{0}a_{1}a_{2}\cdots=0100101100\cdots$ of which all the $B_{i}$, $i\geq 2$, are prefixes.
Define $\beta_{i}=|B_{i}|$, the length of the $i$’th iterate. Let $c_{0}=2$, $c_{1}=4$, $c_{2}=6$, $c_{3}=10$, and define $c_{i}=2c_{i-1}-c_{i-4}$ for $i\geq 4$; this is sequence A288465 in the OEIS. Kimberling conjectured that $\beta_{i}=c_{i}$ for all $i\geq 0$. In this note, we prove Kimberling’s conjecture, as well as his conjectures about the related sequences A288463 and A288464. We also find other properties of the sequence $\bf B$} that link it to the infinite Tribonacci word A080843.
Kimberling indexed the sequence $\bf B$} starting at position $1$. However, for using Walnut, it is easier to index starting at position 0, and this is the convention we use in this paper at the beginning. Later on we will have to use his indexing.
2 Kimberling’s conjecture
In this section we prove Kimberling’s conjecture about the lengths of the words $B_{i}$.
Clearly each $1$ in a $B_{i}$ gives rise to $10$ in $B_{i+1}$, which adds $1$ to the length for each $1$ that appears. Similarly, each $00$ in a $B_{i}$ gives rise to $0101$, which adds $2$ to the length for each $00$ that appears. Defining $N_{w}(i)$ to be the number of occurrences of a word $w$ in $B_{i}$, we therefore have
$\beta_{i+1}=\beta_{i}+N_{1}(i)+2N_{00}(i).$ (1)
Two iterations of the inflation rules turn $11$ into $1001001$, $101$ into $100101100$, and $1001$ into $10010110010100$. Hence
$N_{00}(i+1)=N_{1}(i-1)+N_{00}(i-1).$ (2)
On the other hand the rule $00\rightarrow 0101$ gives
$N_{1}(i+1)=N_{1}(i)+2N_{00}(i).$ (3)
Let $T_{n}$ be the $n$’th Tribonacci number, defined by $T_{0}=0$, $T_{1}=1$, $T_{2}=1$, and $T_{n}=T_{n-1}+T_{n-2}+T_{n-3}$ for $n\geq 3$.
Proposition 1.
We have
$\beta_{i}$ $=T_{i+2}+T_{i}+1$
$N_{00}(i)$ $=T_{i-1}+T_{i-2}$
$N_{1}(i)$ $=2T_{i}$
for $i\geq 2$.
Proof.
By induction on ii. The base cases are easily checked. Now assume the result is true for $i$ and we prove it for $i+1$.
From Eq. (1) we have
$\beta_{i+1}$ $=\beta_{i}+N_{1}(i)+2N_{00}(i)$
$=T_{i+2}+T_{i}+1+2T_{i}+2(T_{i-1}+T_{i-2})$
$=T_{i+3}+T_{i+1}+1.$
From Eq. (2) we have
$N_{00}(i+1)=N_{1}(i-1)+N_{00}(i-1)=2T_{i-1}+T_{i-2}+T_{i-3}=T_{i}+T_{i-1}.$
From Eq. (3) we have
$N_{1}(i+1)=N_{1}(i)+2N_{00}(i)=2T_{i}+2T_{i-1}+2T_{i-2}=2T_{i+1}.$
The proof is complete. ∎
3 Relationship to the Tribonacci word
The infinite sequence $\bf B$ given in A288462 is quite closely related to the celebrated Tribonacci word ${\bf TR}=t_{0}t_{1}t_{2}\cdots=01020100102010102010010201020100102010102010\cdots$, the fixed point of the morphism $0\rightarrow 01$, $1\rightarrow 02$, $2\rightarrow 0$. For more information about $\bf TR$, see, for example, [3].
We will need some additional concepts. Fix an infinite word $\bf x$. We say $w$ is a return word to $y$ in $\bf x$ if ${\bf x}[i..i+n-1]$ and ${\bf x}[j..j+n-1]$ for $i<j$ are two consecutive occurrences of $y$ in $\bf x$ and $w={\bf x}[i..j-1]$. If $y$ occurs with bounded gaps in $\bf x$, we can write $\bf x$ as a concatenation of a finite prefix $v$ and the $t$ different return words to $y$, and hence write ${\bf x}=v\pi({\bf z})$ for some morphism $\pi$ and $\bf z$ a word over $\{0,1,\ldots,t-1\}$. Then $\bf z$ is called the derived sequence of $y$ in $\bf x$ and is denoted by ${\bf d}_{\bf x}(y)$.
Theorem 2.
Kimberling’s sequence $\bf B$} is equal to $0f(\bf TR)$, where $f:0\to 10,\ 1\to 0,\ 2\to 1$.
Proof.
It is easy to see that the return words to $10$ in $\bf B$ are 100, 101, 10100,\ 101,\ 10. If we code $100$ with the letter 0, $101$ with the letter $1$ and $10$ with the letter $2$, we get the derived sequence
${\bf d}_{\bf B}(10)=0102010010201010201001020102\cdots$
[…]
One can check that $\pi=f\circ\varphi$. Hence, ${\bf B}=0\pi({\bf TR})=0f\bigl(\varphi({\bf TR})\bigr)=0f(\bf TR)$ as stated. [^2] We thank to Pascal Ochem who pointed out to us that the morphism $\pi$ can be replaced by the simple morphism $f$.
Our next goal is to find a finite automaton that computes the sequence $\bf B$. For this, we need the notion of Tribonacci representation of an integer.
By a well-known theorem [2], every integer $n\geq 0$ can be written uniquely as a sum of distinct Tribonacci numbers $T_{i}$ for $i\geq 2$, provided one never uses three consecutive Tribonacci numbers in the representation. If we write $n=\sum_{2\leq i\leq t}e_{i}T_{i}$, we can alternatively represent $n$ by the binary word $e_{t}e_{t-1}\cdots e_{2}$. For example, $17=13+4=T_{6}+T_{4}$, and its Tribonacci representation is therefore $10100$.
Our automaton that computes $\bf B$ is a DFAO (deterministic finite automaton with output) computing ${\bf B}[n]$. The input is $n$, expressed in Tribonacci representation, and the output is ${\bf B}[n]$. See, for example, [11].
Let $c_{i}(n)$ denote the number of occurrences of the letter $i$ in the length-$n$ prefix of $\bf TR$. In [15, §10.12], it is shown how to obtain synchronized automata computing the maps $n\rightarrow c_{i}(n)$. By “synchronized” we mean that these automata (called c0, c1, c2) take two inputs in parallel, $n$ and $x$, in Tribonacci representation and accept if and only if $x=c_{i}(n)$. See [14] for more aboout the notion of synchronized automata.
From the description in Theorem 2 that ${\bf B}=0\pi({\bf TR})$, we therefore get the following algorithm for computing ${\bf B}[n]$:
  • If $n=0$ then ${\bf B}[n]=0$.
  • Otherwise, find $y$ such that $3c_{0}(y)+3c_{1}(y)+2c_{2}(y)+1\leq m<3c_{0}(y+1)+3c_{1}(y+1)+2c_{2}(y+1)+1$. This is the position of $\bf TR$ that gives rise to the $n$’th letter of $\bf B$ under the map $\pi$.
  • Let $t=n-(3c_{0}(y)+3c_{1}(y)+2c_{2}(y)+1)$. This is the relative position within the image of ${\bf TR}[y]$ under $\pi$. If $t=0$ it is the first letter, if $t=1$ it is the second letter, and so forth.
  • Then ${\bf B}[n]=1$ if $t=0$, or if $t=2$ and ${\bf TR}[y]=1$; otherwise ${\bf B}[n]=0$.
We can now write Walnut code that implements this algorithm. For more about Walnut and its use in combinatorics on words, see [15]. The first 7 lines are taken from [15, §10.12].
reg shift {0,1} {0,1} "([0,0]|[0,1][1,1]*[1,0])*":
def triba "?msd_trib (s=0&n=0) | Ex \$shift(n-1,x) & s=x+1":
# position of n’th 0 in Tribonacci, starting at index 1
def tribb "?msd_trib (s=0&n=0) | Ex,y \$shift(n-1,x) &
   \$shift(x,y) & s=y+2":
# position of n’th 1 in Tribonacci, starting at index 1
def tribc "?msd_trib (s=0&n=0) | Ex,y,z \$shift(n-1,x) &
   \$shift(x,y) & \$shift(y,z) & s=z+4":
# position of n’th 2 in Tribonacci, starting at index 1
def c0 "?msd_trib Et,u \$triba(s,t) & \$triba(s+1,u) & t<=n & n<u":
def c1 "?msd_trib Et,u \$tribb(s,t) & \$tribb(s+1,u) & t<=n & n<u":
def c2 "?msd_trib Et,u \$tribc(s,t) & \$tribc(s+1,u) & t<=n & n<u":
def find_t_and_y "?msd_trib Eu,v,a0,a1,a2,b0,b1,b2 \$c0(y,a0) & \$c1(y,a1) &
   \$c2(y,a2) & \$c0(y+1,b0) & \$c1(y+1,b1) & \$c2(y+1,b2) & u=3*a0+3*a1+2*a2+1 &
   v=3*b0+3*b1+2*b2+1 & u<=n & n<v & t+u=n"::
# 26 states
def bb "?msd_trib (Et,y \$find_t_and_y(n,t,y) & ((t=0) | (t=2 & TR[y]=@1)))":
combine B bb:
This gives us the automaton in Figure 1.
Now that we have the automaton for $\bf B$, we can use Walnut to provide rigorous proofs of assertions about the sequence $\bf B$. We only need to phrase our assertions in first-order logic, and Walnut can decide if they are TRUE or FALSE. As an example of the utility of the automaton for $\bf B$, we now use Walnut to prove a result about the balance of $\bf B$. A sequence over $\{0,1\}$ is said to be $k$-balanced if for all factors $x,y$ of the same length, the number of $1$’s in $x$ differs from the number of $1$’s in $y$ by at most $k$ [1].
Theorem 3.
The sequence $\bf B$ is $3$-balanced but not $2$-balanced.
Proof.
We can prove this with Walnut. We need an automaton computing bpref1, the number of $1$’s in ${\bf B}[0..n-1]$. We can compute this using the same technique that we used to construct the automaton $\tt B$.
def bpref1 "?msd_trib (n<=1 & z=0) | Et,y,x,a0,a1,a2  \$find_t_and_y(n-1,t,y) &
   \$c0(y,a0) & \$c1(y,a1) & \$c2(y,a2) & z=a0+2*a1+a2+x+1 &
   x<=1 & (x=1 <=> (t=2 & B[n-1]=@1))":
# z = the number of 1’s in B[0..n-1]
def bfact1 "?msd_trib Ex,y \$bpref1(i,x) & \$bpref1(i+n,y) & z+x=y":
# z = the number of 1’s in B[i..i+n-1]
eval bal3 "?msd_trib An,i,j,x,y (\$bfact1(i,n,x) & \$bfact1(j,n,y) & x<=y)
   => y<=x+3":
eval bal2 "?msd_trib An,i,j,x,y (\$bfact1(i,n,x) & \$bfact1(j,n,y) & x<=y)
   => y<=x+2":
The first returns TRUE and the second FALSE. (It fails at $n=47$, as observed by Pierre Popoli.) ∎
Figure 1: The Tribonacci automaton for ${\bf B}[n]$.
4 Related sequences
Kimberling also proposed the sequence A288464, which consists of $I_{1}(n)$, the index of the $n$’th 11 in the sequence $\bf B$, for $n\geq 1$. However, he indexed $\bf B$ starting with $1$. We keep his indexing here. We can compute this with Walnut as follows:
def nth1 "?msd_trib \$bpref1(x,n) & \$bpref1(x-1,n-1)":
# 69 states
Similarly, we can compute $I_{0}(n)$, the index of the $n$’th 0 in $\bf B$, again with $\bf B$ indexed starting at $1$.
def bpref0 "?msd_trib Ey \$bpref1(n,y) & z+y=n":
def nth0 "?msd_trib \$bpref0(x,n) & \$bpref0(x-1,n-1)":
# 30 states
Kimberling conjectured that
$1<\psi-I_{0}(n)/n<1$
for $n\geq 1$ and some constant $\psi\doteq 1.83$. It turns out that $\psi=1.8392867552\cdots$ is the unique real zero of the polynomial $X^{3}-X^{2}-X-1$.
Kimberling also conjectured that
$1<\gamma-I_{1}(n)/n<1$
for $n\geq 1$ and some constant $\gamma\doteq 2.19$. It turns out that $\gamma=(\psi^{2}+1)/2=2.19148788\cdots$.
We will now prove more precise versions of these claims.
Theorem 4.
For $n\geq 1$ we have
  • (a)
    $\lfloor\psi n\rfloor-2\leq I_{0}(n)\leq\lfloor\psi n\rfloor+2$;
  • (b)
    $\lfloor\gamma n\rfloor-1\leq I_{1}(n)\leq\lfloor\gamma n\rfloor+2$.
Proof.
  • (a)
    We use an estimate from [4, Eq. (30)]; namely
    $\lfloor\psi n\rfloor-1\leq A_{0}(n)\leq\lfloor\psi n\rfloor+1$ (4)
    for $n\geq 1$, where $A_{0}(n)$ is the position of the $n$’th 0 in $\bf TR$, where $\bf TR$ is also indexed starting at position $1$. (Similar estimates can be found in [12].)
    We also use some Walnut code from [15] for $A_{0}(n)$, namely the automaton triba.
    Now we show that $1\leq\lfloor\psi n\rfloor-I_{0}(n)\leq 1$:
    eval cmp "?msd_trib An,x,y (\$triba(n,x) & \$nth0(n,y)) =>
       (x=y+1|y=x+1|x=y)":
    And Walnut returns TRUE.
    Putting this together with Eq. (4), we get the estimate
    $2\leq\lfloor\psi n\rfloor-I_{0}(n)\leq 2,$
    from which Kimberling’s first inequality follows easily.
  • (b)
    Let $A_{1}(n)$ denote the position of the $n$’th occurrence of $1$ in the Tribonacci word $\bf TR$ (indexed starting at $1$). In [4, Eq. (31)] the authors showed $\lfloor\psi^{2}n\rfloor-2\leq A_{1}(n)\leq\lfloor\psi^{2}n\rfloor+1$ from which we get
    $1\leq{{A_{1}(n)-\lfloor\psi^{2}n\rfloor}\over 2}\leq 1/2$ (5)
    by rearranging.
    On the other hand, we can use Walnut to prove that
    $A_{1}(n)\leq 2I_{1}(n)+1-n\leq A_{1}(n)+5.$
    For $A_{1}(n)$ we use the code tribb from [15, §10.12]:
    eval comp2 "?msd_trib An,x,y,z (\$tribb(n,x) & \$nth1(n,y) & z+n=2*y+1)
       => (x<=z & z<=x+5)":
    And Walnut returns TRUE. From this we get
    $0\leq{{2I_{1}(n)-n+1-A_{1}(n)}\over 2}\leq 5/2.$
    Adding this to Eq. (5) we get
    $1\leq{{2I_{1}(n)-n-\lfloor\psi^{2}n\rfloor+1}\over 2}\leq 3$
    which is equivalent to
    $3/2\leq I_{1}(n)-\lfloor\gamma n\rfloor\leq 5/2$
    for $\gamma=(\psi^{2}+1)/2$. Since $I_{1}(n)-\lfloor\gamma n\rfloor$ is an integer, we get
    $1\leq I_{1}(n)-\lfloor\gamma n\rfloor\leq 2,$
    from which Kimberling’s second inequality now follows easily.
5 Subword complexity and critical exponent
We use the term “subword” and “factor” to mean the same concept: a contiguous block of letters within a word.
Recall that the subword complexity (aka factor complexity) of a sequence is the function mapping $n$ to the number of distinct blocks of length $n$ appearing in it. A word is called a Rote word [13] if its subword complexity is $2n$ for $n\geq 1$.
We will also need the notation of critical exponent of an infinite word. We say a finite word $w=w[1..n]$ has period $p$ if $w[i]=w[i+p]$ for all $i$, $1\leq i\leq n-p$. The smallest nonzero period is called the period and is denoted $\operatorname{per}(w)$. The exponent of a finite nonempty word $w$ is then $\exp(w):=|w|/\operatorname{per}(w)$. Let $\bf x$ be an infinite word. Then the critical exponent of $\bf x$, written $\operatorname{ce}({\bf x})$, is $\sup\{\exp(w)\,:\,w\text{ is a factor of {\bf x}}\}$.
The goal of this section is to prove that the sequence $\bf B$ belongs to the class of Rote words and has the same critical exponent as that of the Tribonacci word [16].
More specifically, we aim to prove the following two theorems.
Theorem 5.
The subword complexity of $\bf B$ is $2n$ for $n\geq 1$.
Theorem 6.
The critical exponent of $\bf B$ is $2+\frac{1}{\psi-1}=3.19148788395\cdots$, where $\psi$ is the real zero of $X^{3}-X^{2}-X-1$.
In principle, Walnut could be used to prove both of these theorems; but in practice, we were unable to complete the proof because the computations fail to terminate within reasonable bounds on space and time. So in the next section, we use some known theory instead.
5.1 Bispecial factors of the sequence $\bf B$
For a binary sequence $\bf x$, we say a factor $w$ is right-special if $w0$ and $w1$ both appear in $\bf x$, and left-special if both $0w$ and $1w$ both appear in $\bf x$. If $w$ is both right- and left-special, we say it is bispecial.
In order to determine both the subword complexity and the critical exponent of $\bf B$, the knowledge of bispecial factors in $\bf B$ is essential. The description of bispecial factors and their return words in $\bf TR$ is taken from [6, 8, 7]. The sequence $(b_{n})_{n=0}^{\infty}$ of all non-empty bispecial factors, ordered by length, in the Tribonacci word, satisfies $b_{0}=0$ and for $n\geq 1$,
$b_{n}=\varphi(b_{n-1})0,\quad\text{where}\ \varphi:0\to 01,\ 1\to 02,\ 2\to 0\,.$
Moreover, if $i\equiv n\ ({\rm mod}\ 3)$ for $i\in\{-1,0,1\}$, then the both-sided extensions of $b_{n}$ are
$(i+1)b_{n}j,\ jb_{n}(i+1)\quad\text{for}\ j\in\{0,1,2\}\,.$ (6)
By Theorem 2, Kimberling’sequence is equal to $f({\bf TR})$. As $\bf TR$} is fixed by $\varphi$, $\bf B$} is also the image of the Tribonacci sequence under the morphism $\pi=f\circ\varphi$. Instead of $f$ we will work with the morphism $\pi:0\to 100,\ 1\to 101,\ 2\to 10$ because it allows us to derive the form of bispecial factors in $\bf B$ in an easier way.
Observation 7.
The non-empty bispecial factors of $\bf B$ of length $\leq 4$ are $0,1,01,10,010$.
Lemma 8.
The complete list of bispecial factors of $\bf B$ of length $\geq 5$ is as follows:
  1. 1.
    for $n\equiv 0\ ({\rm mod}\ 3)$:
    $\pi(b_{n})10,\ \pi(b_{n})101\,;$
  2. 2.
    for $n\equiv 1\ ({\rm mod}\ 3)$:
    $\pi(b_{n})10,\ \pi(b_{n})101,\ 0\pi(b_{n})10,\ 0\pi(b_{n})101\,;$
  3. 3.
    for $n\equiv 2\ ({\rm mod}\ 3)$:
[…]
Proof.
[…]
$\pi(1b_{n}0)$ $=101\pi(b_{n})100,$
$\pi(0b_{n}10)$ $=100\pi(b_{n})101100,$
$\pi(1b_{n}20)$ $=101\pi(b_{n})10100$
are factors of $\bf B$, it follows that $\pi(b_{n})10$ and $\pi(b_{n})101$ are bispecial factors in $\bf B$. On the other hand, each bispecial factor $w$ in $\bf B$ of length at least 5 starts with $10$ or $010$ and ends with $10$ or $101$. By the form of the morphism $\pi$, the factor $w$ takes one of the following forms
$w\in\{\pi(b)10,\ \pi(b)101,\ 0\pi(b)10,\ 0\pi(b)101\}\,$,
where $b$ is a non-empty bispecial factor in $\bf TR$. Consequently, the factor $w$ is included in the list from Lemma 8. ∎
5.2 Subword complexity of the sequence $\bf B$
Lemma 9.
The set of left special factors of $\bf B$ is equal to the set of prefixes of $\pi(\bf TR)$ and of $\bf B=0\pi(\bf TR)$.
Proof.
Since $\bf B$ is aperiodic, each left special factor is the prefix of a bispecial factor. All bispecial factors of $\bf B$ are prefixes of $\bf B$ or $\pi(\bf TR)$. This statement is clear for bispecial factors of length $\leq 4$. It remains to see that the bispecial factors of length $\geq 5$, as listed in Lemma 8, are prefixes of $\bf B$ or $\pi(\bf TR)$. It can be easily proven by induction that $b_{n}(i+1)$ is the prefix of $\bf TR$ for $n\equiv i\ ({\rm mod}\ 3)$, where $i\in\{-1,0,1\}$. Therefore, for instance for $n\equiv 0\ ({\rm mod}\ 3)$, the bispecial factors $\pi(b_{n})10$ and $\pi(b_{n})101$ are prefixes of $\pi(b_{n}1)$, which is a prefix of $\pi(\bf TR)$. We can proceed analogously for $n\equiv 1\ ({\rm mod}\ 3)$ and $n\equiv 2\ ({\rm mod}\ 3)$. ∎
Proof of Theorem 5.
By Lemma 9, for each length $n\geq 1$, there are two distinct left special factors in $\bf B$, which confirms that the subword complexity is $2n$ for all $n\geq 1$. ∎
5.3 Return words to bispecial factors in $\bf B$
For the purpose of computing the critical exponent of $\bf B$, we intend to apply the following theorem.
Theorem 10 ([5], Theorem 3).
Let $\bf u$ be a uniformly recurrent aperiodic sequence. Let $(w_{n})_{n\in\mathbb{N}}$ be the sequence of all bispecial factors in $\bf u$ ordered by length. For every $n\in\mathbb{N}$, let $v_{n}$ be the shortest return word to the bispecial factor $w_{n}$ in $\bf u$. Then
$\operatorname{ce}({\bf u})=1+\sup\left\{\frac{|w_{n}|}{|v_{n}|}\ :\ n\in\mathbb{N}\right\}\,.$
It is thus necessary to describe the shortest return words to bispecial factors in $\bf B$.
There are three return words to each factor in the Tribonacci word. In particular, the return words to $b_{n}$, for $n\geq 0$, are
$\varphi^{n}(0),\ \varphi^{n}(01),\ \varphi^{n}(02)\,.$
Proposition 11.
Let $w$ be a bispecial factor from the list in Lemma 8.
  • (a)
    If $w=\pi(b_{n})10$, then the shortest return word to $w$ equals $\pi(\varphi^{n}(0))$;
  • (b)
    If $w=\pi(b_{n})101$ or $b=0\pi(b_{n})10$, then each return word to $w$ has length $\geq|\pi(\varphi^{n}(0))|$;
  • (c)
    If $w=0\pi(b_{n})101$, then each return word to $w$ has length $\geq|\pi(\varphi^{n}(0))|+|\pi(\varphi^{n-1}(0))|$.
Proof.
Using the form of $\pi$ and its injectivity, the factor $\pi(b_{n})$ has a unique preimage $b_{n}$, therefore the shortest complete return word $\varphi^{n}(0)b_{n}$ to $b_{n}$ gives rise to the shortest complete return word $\pi(\varphi^{n}(0)b_{n})$ to $\pi(b_{n})$.
  • (a)
    Since $\pi(b_{n})$ is always followed by $10$, the factor $\pi(b_{n})10$ has the same shortest return word as $\pi(b_{n})$. This proves (a).
  • (b)
    The claim (b) immediately follows from (a).
  • (c)
    By Lemma 8, the factor $w=0\pi(b_{n})101$ is bispecial only for $n\equiv 1\ ({\rm mod}\ 3)$. Since the last letter of $\varphi^{n}(0)$ equals $i$, where $n\equiv i\ ({\rm mod}\ 3)$, the factor $0\pi(\varphi^{n}(0)b_{n})101$ has the suffix $101\pi(b_{n})101$, which proves that $0\pi(\varphi^{n}(0))0^{-1}$ is not a return word to $0\pi(b_{n})101$.
    Since $0\pi(b_{n})101$ contains $\pi(b_{n})$, for each of its return words $v$, the word $0^{-1}v0$ is obtained as a concatenation of return words to $\pi(b_{n})$. This concatenation is not equal to $\pi(\varphi^{n}(0))$, and hence
    $|v|\geq\min\{|\pi(\varphi^{n}(00))|,|\pi(\varphi^{n}(01))|,|\pi(\varphi^{n}(02))|\}=|\pi(\varphi^{n}(02))|=|\pi(\varphi^{n}(0))|+|\pi(\varphi^{n-1}(0))|.$
5.4 Critical exponent of the sequence $\bf B$
In order to apply Theorem 10, we need to determine the lengths of bispecial factors and their shortest return words in $\bf B$.
Recall that in $\bf TR$, the sequence $(b_{n})_{n=0}^{\infty}$ of all non-empty bispecial factors satisfies $b_{0}=0$ and $b_{n}=\varphi(b_{n-1})0$ for $n\geq 1$ and $r_{n}=\varphi^{n}(0)$ is the shortest return word to $b_{n}$. Also recall that the Parikh vector of a word $x$ over the alphabet $\{0,1,\ldots,t-1\}$ is the vector $(|x|_{0},|x|_{1},\ldots,|x|_{t-1}),$ where $|x|_{a}$ is the number of occurrences of $a$ in $x$.
The Parikh vectors of bispecial factors and their shortest return words in $\bf TR$ are
$\vec{b}_{n}=\frac{1}{2}\left(\begin{array}[]{c}T_{n+3}+T_{n+1}-1\\ T_{n+2}+T_{n\phantom{-1}}-1\\ T_{n+1}+T_{n-1}-1\end{array}\right)\quad\text{and}\quad\vec{r}_{n}=\left(\begin{array}[]{c}T_{n+1}\\ T_{n\phantom{-1}}\\ T_{n-1}\end{array}\right).$ (7)
The explicit form of $T_{n}$ reads, for $n\geq 0$,
$T_{n}=c_{1}\psi_{1}^{n}+c_{2}\psi_{2}^{n}+c_{3}\psi_{3}^{n}\,$, (8)
where $\psi_{1}=\psi\doteq 1.8393,\ \psi_{2}=\overline{\psi_{3}}\doteq-0.4196+0.6063i$, and $c_{j}=\frac{1}{-\psi_{j}^{2}+4\psi_{j}-1}$ for $j\in\{1,2,3\}$.
The following lemma enables to express the lengths of all bispecial factors and their shortest return words in $\bf B$ in terms of the Tribonacci numbers.
Lemma 12.
For $n\geq 0$ we have
$|\pi(r_{n})|=T_{n+5}-T_{n+4}\quad\text{and}\quad|\pi(b_{n})|=|\pi(r_{n})|+T_{n+4}-4\,.$
Proof.
Using the Tribonacci recurrence, we get
$|\pi(r_{n})|=(1,1)M_{\pi}\vec{r}_{n}=(1,1)\left(\begin{array}[]{rcl}2&1&1\\ 1&2&1\end{array}\right)\left(\begin{array}[]{c}T_{n+1}\\ T_{n\phantom{-1}}\\ T_{n-1}\end{array}\right)=T_{n+5}-T_{n+4}\,.$
$|\pi(b_{n})|=(1,1)M_{\pi}\vec{b}_{n}=(1,1)\left(\begin{array}[]{rcl}2&1&1\\ 1&2&1\end{array}\right)\frac{1}{2}\left(\begin{array}[]{c}T_{n+3}+T_{n+1}-1\\ T_{n+2}+T_{n\phantom{-1}}-1\\ T_{n+1}+T_{n-1}-1\end{array}\right)=T_{n+5}-4\,.$
Proof of Theorem 6.
Combining Theorem 10, Lemma 8, Proposition 11, and Lemma 12, we have
$\operatorname{ce}({\bf B})\geq 1+\lim_{n\to\infty}\frac{|\pi(b_{n})10|}{|\pi(r_{n})|}=2+\lim_{n\to\infty}\frac{T_{n+4}-2}{T_{n+5}-T_{n+4}}=2+\frac{1}{\psi-1}.$
Now, for every bispecial factor $w$ of length $\leq 4$ in $\bf B$ and its shortest return word $v$, the inequality $1+\frac{|w|}{|v|}\leq 2+\frac{1}{\psi-1}\doteq 3.19$ holds. Here is a table of such bispecial factors and their shortest return words.
$\begin{array}[]{c|c|c|c|c|c}w&0&1&01&10&010\\ \hline\cr v&0&1&01&10&01\\ \hline\cr|w|/|v|&1&1&1&1&1.5\end{array}$
To complete the proof that $\operatorname{ce}({\bf B})\leq 2+\frac{1}{\psi-1}=1+\frac{\psi}{\psi-1}$, it suffices to show that for all $n\geq 0$ we have
$\frac{|\pi(b_{n})101|}{|\pi(r_{n})|}\leq\frac{\psi}{\psi-1}\quad\text{and}\quad\frac{|0\pi(b_{n})101|}{|\pi(r_{n})|+|\pi(r_{n-1})|}\leq\frac{\psi}{\psi-1}\,.$ (9)
The first inequality from (9) may be simplified as follows:
$\frac{T_{n+5}-1}{T_{n+5}-T_{n+4}}\leq\frac{\psi}{\psi-1}\,,\quad\text{or equivalently }\quad\psi\leq\frac{T_{n+5}-1}{T_{n+4}-1}\,.$ (10)
The second inequality from (9) can be rewritten as
$\frac{T_{n+5}}{T_{n+5}-T_{n+3}}\leq\frac{\psi}{\psi-1}\,,\quad\text{or equivalently }\quad\psi\leq\frac{T_{n+5}}{T_{n+3}}\,.$
Since $\frac{T_{n+5}}{T_{n+3}}\geq\frac{T_{n+5}-1}{T_{n+3}}\geq\frac{T_{n+5}-1}{T_{n+4}-1}$, only the first inequality (10) needs to be verified. It obviously holds for $n\in\{0,1\}$. Using the explicit formula (8) for $T_{n}$, we see that for $n\geq 2$ we have
$T_{n}-c_{1}\psi^{n}\in(-K,K),\ \ \text{where }K\leq 2|c_{2}\psi^{2}_{2}|.$
For $\psi\doteq 1.8393$, $c_{2}\doteq-0.1681+0.1983i$ and $\psi_{2}\doteq-0.4196+0.6063i$, the parameter $K$ satisfies $K\leq 0.29$, and the inequality $\psi\geq\frac{K+1}{1-K}$ holds. Hence
$\frac{T_{n+5}-1}{T_{n+4}-1}\geq\frac{c_{1}\psi^{n+5}-1-K}{c_{1}\psi^{n+4}-1+K}\geq\psi.$
The proof is now complete. ∎
6 Acknowledgments
We acknowledge with thanks conversations with Pierre Popoli.